# A Scoping Review of Natural Language Processing in Addressing Medically Inaccurate Information: Errors, Misinformation, and Hallucination

### v1

@@ Abstract
Methods: A scoping review was conducted following PRISMA guidelines, analyzing studies from 2020 to 2024 across five databases. Studies were selected based on their use of NLP to address medically inaccurate information and were categorized by topic, tasks, document types, datasets, models, and evaluation metrics.
@@ 1 Introduction
Advancements in Natural Language Processing (NLP) and LLMs are revolutionizing the way we detect and correct medically inaccurate information. Transformer-based models like Bidirectional Encoder Representations from Transformers (BERT) [20] and their medical domain adaptations, such as BioBERT [21] and ClinicalBERT [22], have been utilized for error detection and classification in clinical texts [23]. NLP techniques also played a critical role in misinformation detection by monitoring social media and online forums for health-related inaccuracies. During the COVID-19 pandemic, models were trained to detect COVID-19-related misinformation from tweets by classifying content based on veracity [24]. Additionally, tasks such as fact-checking and claim verification leveraged NLP combined with knowledge graphs to verify claims against trusted medical databases, including PubMed and guidelines from the Centers for Disease Control and Prevention (CDC) [25, 26]. LLMs such as GPT-4 [27] exhibited unique strengths in error correction by generating contextually rich, human-like explanations and reasoning through interactive dialogues [28]. These capabilities, supported by advancements in natural language understanding (NLU), enabled LLMs to integrate domain-specific knowledge and address complex inaccuracies with greater flexibility [29, 30].
@@ 2 Methods
Our scoping review follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines [42].
@@ 2.4 Study Selection
Both GPT-4o and a human reviewer (ZS) independently conducted title and abstract screenings. The Cohen’s kappa was 0.70, indicating substantial agreement between GPT-4o and ZS. Besides the exclusion criteria mentioned in the prompts above, papers were also excluded if they were inaccessible or published prior to 2020. If ZS and GPT-4o disagreed on a paper, ZS conducted a double-check and decided whether to include it in the next stage. To evaluate the reliability of ZS’s decisions, a random sample of 40 papers was independently assessed by two additional reviewers (OU and MY), and all decisions were consistent with those made by ZS. Papers selected during this screening process then moved to a full-text review by ZS. All included papers and their categorization were discussed with all co-authors before finalizing selections.
@@ 3 Results
Figure 2 shows the flowchart of article selection, following PRISMA guidelines. A total of 1,543 articles were initially identified from five databases. Of these, 542 articles were excluded prior to screening for reasons such as duplication, non-research content, publication before 2020, or access issues, leaving 1,001 articles for a quick title screening. Following this initial screening, 589 articles were excluded, and 412 articles proceeded to a detailed title and abstract screening conducted by the human reviewer, assisted by GPT-4o. At this stage, 174 additional articles were excluded based on the criteria outlined in Section 2.4. Subsequently, 238 articles proceeded to full-text review, during which an additional 183 articles were excluded: 8 were non-medical, 19 had overly broad scopes, 6 lacked a focus on NLP, 22 contained no inaccurate information, and 16 were of poor quality. Additionally, 112 articles were excluded due to overlap in topics or methodologies. For instance, numerous studies on COVID-19-related misinformation detection published between 2020 and 2022 employed similar study designs and models. To avoid redundancy, only a representative selection of these studies was included. Similarly, shared tasks in the field often result in multiple papers addressing the same topic, all meeting our inclusion criteria. In such cases, we included only a few top-ranked or highly innovative papers. Ultimately, 55 articles were included in our final review, categorized by type of inaccuracy: 13 focused on errors, 27 on misinformation, and 15 on hallucination.
@@
| MEDIQA-CORR 2024 [45] | general medical | 1 | medical question-answering text and clinical notes with injected errors | English | text | binary | https://github.com/abachaa/MEDIQA-CORR-2024 |
@@ 4 Errors
| Valiev et al. [77] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDIQA-CORR 2024 | GPT-3.5, GPT-4 | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
| Toma et al. [29] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDIQA-CORR 2024 | GPT-3.5, GPT-4, DSPy framework | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
| Gundabathula et al. [28] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDIQA-CORR 2024 | GPT-3.5, GPT-4, Claude-3 Opus | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
@@ 4.1 Error detection
Multi-label classification tasks assign multiple error categories to a single instance. Wong et al. [73] applied deep neural network (DNN) models (a feedforward artificial neural network model with 1-5 hidden layers) to classify medication administration errors, focusing on wrong patient, wrong drug, wrong time, wrong dose, and wrong route. The dataset consisted of 574 incident reports collected from the Hong Kong Hospital Authority’s Advanced Incident Reporting System (AIRS) over 2011–2014, with structured and unstructured free-text fields describing medication errors and near-misses. DNN models achieved high performance, with an average accuracy of 0.94 and an AUC of 0.911 across all five categories. Boxley et al. [75] employed logistic regression, elastic net, and XGBoost models to classify medication errors in patient safety event (PSE) reports using eight categories adapted from the NCC MERP taxonomy [82], such as wrong drug, wrong time, wrong dose, and monitoring errors. The dataset included 3,861 annotated PSE reports from a ten-hospital healthcare system, with structured fields and free-text narratives describing medication safety events. Among the models tested, XGBoost demonstrated the highest performance, with an average F1-score of 0.72 across categories. These error types and categories play a crucial role in addressing inaccurate information by identifying underlying patterns of medication errors, ultimately enhancing safety processes and reducing the likelihood of similar mistakes in the future.
LLM-based classification methods integrate error detection into the inference process to enhance downstream tasks. Gundabathula et al. [28] categorized errors into domains (e.g., medications, medical conditions, clinical procedures) and included this structure in model prompts. This guided the model through a CoT reasoning process, which improved the model’s accuracy and ensured more precise and explainable results. It also helped reduce hallucination and enhance consistency.
@@ 4.2 Error correction
In the reviewed articles, error correction primarily includes spelling correction and contextual error correction. Spelling correction focuses on typographical and lexical errors, while contextual error correction addresses issues such as incorrect diagnoses, treatments, or medication instructions derived from broader contextual information. In spelling correction tasks, the most common evaluation metrics are precision, recall, and F1. Bravo-Candel et al. [44] used F0.5 instead of F1, giving more weight to precision. This choice was made because in spelling correction tasks, false positives are often undesirable, making models with higher precision (fewer false positives) preferable. In contextual error detection tasks, accuracy was used to evaluate the correct identification of errors (error flagging) and the correct detection of sentences containing errors [83]. In contextual error correction tasks, ROUGE, BERTsocre, and BLEURT were commonly used. ROUGE measured unigram overlap between generated and reference text, commonly used for sentence correction [84]. BERTScore used contextual embeddings to assess semantic similarity between generated and reference sentences [85], while BLEURT employed machine-learned metrics trained on human ratings for deeper quality evaluation [86]. AggregateScore combined multiple metrics to provide a balanced measure of performance across different dimensions, and was used for ranking in MEDIQA-CORR 2024 [83]. Only one paper involved clinical expert evaluating the system’s outputs to identify potential near-miss events and validate the safety and accuracy of medication directions [78].
@@
Two papers specifically focused on spelling correction. Bravo-Candel et al. [44] used Seq2Seq neural machine translation models to correct real-word errors in Spanish clinical texts. The study used two datasets: the Wikicorpus, comprising over 611 million words from Spanish Wikipedia articles, and the medicine corpus, a smaller dataset of approximately 5,750 clinical cases with around 2 million words from three Spanish clinical corpora (CodiEsp [87], MEDDOCAN [88], SPACCC [89]). Errors were synthetically introduced in the sentences using predefined rules across six categories, such as grammatical gender and subject-verb concordance. The best performance was achieved with models trained on the medicine corpus, yielding an F0.5 score of 0.6498 with no pre-trained embeddings (GloVe or Word2Vec). The Spanish-language dataset used in this study is one of the few publicly available clinical datasets for error detection and correction. Lee et al. [23] employed a Masked Language Model (MLM)-based approach to correct spelling errors in unstructured medical texts and enhance NER accuracy. The study utilized two datasets: the NCBI-disease corpus [90] (793 PubMed abstracts annotated with disease mentions) and surgical pathology records (40,443 annotated lung cancer diagnostic records from the Asan Medical Center). Synthetic errors were introduced to mimic real-world typographical patterns. The MLM-based spelling correction achieved F1-scores of 0.72 (NCBI-disease) and 0.73 (surgical pathology records). For NER tasks, spelling correction significantly boosted F1-scores in surgical pathology records from 0.60 to 0.85. This study highlighted the potential of spelling correction models to mitigate data quality issues and improve the accuracy of downstream NLP tasks in clinical settings.
@@
Contextual error correction has seen advancements through the application of LLMs for addressing complex, context-dependent inaccuracies. A highlight in this domain is the MEDIC system introduced by Pais et al. [78], which focuses on preventing medication direction errors in online pharmacies by improving accuracy and standardization during the data entry phase. The study utilized 1.6 million single-line medication directions from Amazon Pharmacy, including raw prescriber directions and pharmacist-verified equivalents. MEDIC employs a three-stage process: pharmalexical normalization, AI-powered extraction using a fine-tuned DistilBERT model, and semantic assembly with safety guardrails informed by a medication catalog derived from RxNorm, OpenFDA, and Amazon Pharmacy data. Compared to T5-FineTuned and Claude, MEDIC reduced near-miss events by 33 during deployment, with a notable improvement in suggestion adoption rates and a reduction in post-adoption edits.
@@ 4.3 Others
Tavabi et al. [79] highlighted a critical issue in codification errors related to gold standard labels while evaluating TF-IDF, Doc2Vec, and BERT models for assigning current procedural terminology (CPT) codes from operative notes. The study, focused on musculoskeletal procedures, used a dataset of 44,002 operative notes annotated with the 100 most common CPT codes. TF-IDF demonstrated superior performance with an AUROC of 0.96 and accuracy of 0.97, outperforming both Doc2Vec and BERT, likely due to its robustness to data sparsity and noise in clinical notes. Importantly, the study revealed discrepancies in gold standard CPT assignments during experiments, with NLP models identifying mislabeled instances, thereby outperforming the provided ground truth in some cases. This underscores the need for refining data labeling processes to enhance the reliability of automated codification systems.
@@ 5 Misinformation
Misinformation correction is less explored compared to detection, with only two studies in this review addressing it. Both studies focused on generating polite, evidence-backed responses using social media posts and academic articles as sources of misinformation and supporting evidence. The two papers either released datasets or built on publicly available datasets, but further datasets are still needed to support advancements in this area.
@@ 5.1.2 Fact-checking and claim verification
Validation and interpretation is the final step in fact-checking and claim verification, focusing on model explainability. This stage aims to clarify the reasoning behind predictions and increase trust and transparency in the model’s conclusions. However, this step is often overlooked, with only a few studies incorporating explanation generation and human evaluation. Kotonya et al. [64] used human evaluations of explanations, where annotators assessed coherence and relevance to ensure that the generated explanations were logically consistent with the claim and label. In Tan et al. [62], human annotators evaluated claims generated from multiple-choice questions for qualities like fluency, contextuality, and faithfulness. While some studies included expert evaluations for claim quality or evidence matching, few employed specific metrics to measure interpretability. In future studies, effective validation and interpretation methods will be essential for creating more transparent fact-checking systems that can explain decisions comprehensively and reliably.
@@ 6 Hallucination
| Van Veen et al. [130] | general medical | 1 | LLM-generated summaries | Open-i, MIMIC-CXR, MIMIC-III, MeQSum, ProbSum GPT-4, GPT-3.5, FLAN-T5, FLAN-UL2, Vicuna, Llama-2 | BLEU, ROUGE-L, BERTScore, MEDCON, correctness, completeness, conciseness | automatic, expert | |

### v2

@@ Abstract
Methods: A scoping review was conducted following PRISMA-ScR guidelines, analyzing studies from 2020 to 2024 across five databases. Studies were selected based on their use of NLP to address medically inaccurate information and were categorized by topic, tasks, document types, datasets, models, and evaluation metrics.
@@ 1 Introduction
Advancements in Natural Language Processing (NLP), especially LLMs, are revolutionizing the way we detect and correct medically inaccurate information. Transformer-based models like Bidirectional Encoder Representations from Transformers (BERT) [20] and their medical domain adaptations, such as BioBERT [21] and ClinicalBERT [22], have been utilized for error detection and classification in clinical texts [23]. NLP techniques also played a critical role in misinformation detection by monitoring social media and online forums for health-related inaccuracies. During the COVID-19 pandemic, models were trained to detect COVID-19-related misinformation from tweets by classifying content based on veracity [24]. Additionally, tasks such as fact-checking and claim verification leveraged NLP combined with knowledge graphs to verify claims against trusted medical databases, including PubMed and guidelines from the Centers for Disease Control and Prevention (CDC) [25, 26]. Although encoder-only models like BERT can also be referred to as LLMs [27, 28], we followed the more common usage that restricts the term “LLM” for generative models. LLMs such as GPT-4 [29] exhibited unique strengths in error correction by generating contextually rich, human-like explanations and reasoning through interactive dialogues [30]. These capabilities, supported by advancements in natural language understanding (NLU), enabled LLMs to integrate domain-specific knowledge and address complex inaccuracies with greater flexibility [31, 32].
@@ 2 Methods
Our scoping review follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses Extension for Scoping Reviews (PRISMA-ScR) guidelines [44].
@@ 2.4 Study Selection
Both GPT-4o and a human reviewer (ZS) independently conducted title and abstract screenings. The Cohen’s kappa was 0.70, indicating substantial agreement between GPT-4o and ZS. Besides the exclusion criteria mentioned in the prompts above, papers were also excluded if they were inaccessible or published prior to 2020. If ZS and GPT-4o agreed on a paper, the agreed decision was accepted directly. If ZS and GPT-4o disagreed on a paper, ZS conducted a double-check and decided whether to include it in the next stage. To evaluate the reliability of both agreed and disagreed decisions, a random sample of 40 papers was independently assessed by two additional reviewers (OU and MY), including 10 papers from each of the following agreement/disagreement categories: (1) both included, (2) ZS included but GPT-4o did not, (3) GPT-4o included but ZS did not, and (4) both excluded. All decisions in this sample were consistent with those made by ZS. Papers selected during this screening process then moved to a full-text review by ZS. All included papers and their categorization were discussed with all co-authors before finalizing selections.
@@ 3 Results
Figure 2 shows the flowchart of article selection, following PRISMA-ScR guidelines. A total of 1,543 articles were initially identified from five databases. Of these, 542 articles were excluded prior to screening for reasons such as duplication, non-research content, publication before 2020, or access issues, leaving 1,001 articles for a quick title screening. Following this initial screening, 589 articles were excluded, and 412 articles proceeded to a detailed title and abstract screening conducted by the human reviewer, assisted by GPT-4o. At this stage, 174 additional articles were excluded based on the criteria outlined in Section 2.4. Subsequently, 238 articles proceeded to full-text review, during which an additional 183 articles were excluded: 8 were non-medical, 19 had overly broad scopes, 6 lacked a focus on NLP, 22 contained no inaccurate information, and 16 were of poor quality. Articles with “poor quality” met the initial inclusion criteria but were excluded during full-text review due to issues such as lack of methodological transparency, minimal/vague use of NLP, or absence of NLP evaluation metrics. Additionally, 112 articles were excluded due to overlap in topics or methodologies. For instance, numerous studies on COVID-19-related misinformation detection published between 2020 and 2022 employed similar study designs and models. To avoid redundancy, we reviewed all eligible studies and, when multiple papers shared similar topics, document type, and NLP methods, we randomly selected one to include. Similarly, shared tasks in the field often result in multiple papers addressing the same topic, all meeting our inclusion criteria. In such cases, we included only top 3 ranked submissions. Ultimately, 55 articles were included in our final review, categorized by type of inaccuracy: 13 focused on errors, 27 on misinformation, and 15 on hallucination.
@@
| MEDEC [47] | general medical | 1 | medical question-answering text and clinical notes with injected errors | English | text | binary | https://github.com/abachaa/MEDIQA-CORR-2024 |
@@ 4 Errors
| Valiev et al. [79] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDEC | GPT-3.5, GPT-4 | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
| Toma et al. [31] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDEC | GPT-3.5, GPT-4, DSPy framework | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
| Gundabathula et al. [30] | general medical | 1, 2 | medical question-answering text and clinical notes with injected errors | MEDEC | GPT-3.5, GPT-4, Claude-3 Opus | accuracy, ROUGE, BERTScore, BLEURT, AggregateComposite, AggregateScore | automatic |
@@ 4.1 Error detection
Multi-label classification tasks assign multiple error categories to a single instance. Wong et al. [75] applied deep neural network (DNN) models (feedforward artificial neural networks with varying architectures consisting of 1-5 hidden layers) to classify medication administration errors, focusing on wrong patient, wrong drug, wrong time, wrong dose, and wrong route. The dataset consisted of 574 incident reports collected from the Hong Kong Hospital Authority’s Advanced Incident Reporting System (AIRS) over 2011–2014, with structured and unstructured free-text fields describing medication errors and near-misses. DNN models achieved high performance, with an average accuracy of 0.94 and an AUC of 0.911 across all five categories. Boxley et al. [77] employed logistic regression, elastic net, and XGBoost models to classify medication errors in patient safety event (PSE) reports using eight categories adapted from the NCC MERP taxonomy [84], such as wrong drug, wrong time, wrong dose, and monitoring errors. The dataset included 3,861 annotated PSE reports from a ten-hospital healthcare system, with structured fields and free-text narratives describing medication safety events. Among the models tested, XGBoost demonstrated the highest performance, with an average F1-score of 0.72 across categories. These error types and categories play a crucial role in addressing inaccurate information by identifying underlying patterns of medication errors, ultimately enhancing safety processes and reducing the likelihood of similar mistakes in the future.
LLM-based classification methods integrate error detection into the inference process to enhance downstream tasks. Gundabathula et al. [30] categorized errors into domains (e.g., medications, medical conditions, clinical procedures) and included this structure in model prompts. This guided the model through a CoT reasoning process, which improved the model’s accuracy and ensured more precise and explainable results. In GPT-3.5, classification accuracy increased from 48.75 to 58.44, and span identification accuracy from 22.5 to 38.55. Using GPT-4 further improved performance, reaching 63.07 and 58.17 for those tasks, respectively. It also helped reduce hallucination and enhance consistency.
@@ 4.2 Error correction
In the reviewed articles, error correction primarily includes spelling correction and contextual error correction. Spelling correction focuses on typographical and lexical errors, while contextual error correction addresses issues such as incorrect diagnoses, treatments, or medication instructions derived from broader contextual information. In spelling correction tasks, the most common evaluation metrics are precision, recall, and F1. Bravo-Candel et al. [46] used F0.5 instead of F1, giving more weight to precision. This choice was made because in spelling correction tasks, false positives are often undesirable, making models with higher precision (fewer false positives) preferable. In contextual error detection tasks, accuracy was used to evaluate the correct identification of errors (error flagging) and the correct detection of sentences containing errors [85]. In contextual error correction tasks, ROUGE, BERTscore, and BLEURT were commonly used. ROUGE measured unigram overlap between generated and reference text, commonly used for sentence correction [86]. BERTScore used contextual embeddings to assess semantic similarity between generated and reference sentences [87], while BLEURT employed machine-learned metrics trained on human ratings for deeper quality evaluation [88]. AggregateScore combined multiple metrics to provide a balanced measure of performance across different dimensions, and was used for ranking in MEDIQA-CORR 2024 [85]. Only one paper involved clinical expert evaluating the system’s outputs to identify potential near-miss events and validate the safety and accuracy of medication directions [80].
@@
Two papers specifically focused on spelling correction. Bravo-Candel et al. [46] used Seq2Seq neural machine translation models to correct real-word errors in Spanish clinical texts. The study used two datasets: the Wikicorpus, comprising over 611 million words from Spanish Wikipedia articles, and the medicine corpus, a smaller dataset of approximately 5,750 clinical cases with around 2 million words from three Spanish clinical corpora (CodiEsp [89], MEDDOCAN [90], SPACCC [91]). Errors were synthetically introduced in the sentences using predefined rules across six categories, such as grammatical gender and subject-verb concordance. The best performance was achieved with models trained on the medicine corpus, yielding an F0.5 score of 0.6498 with no pre-trained embeddings. The Spanish-language dataset used in this study is one of the few publicly available clinical datasets for error detection and correction. Lee et al. [23] employed a Masked Language Model (MLM)-based approach to correct spelling errors in unstructured medical texts and enhance NER accuracy. The study utilized two datasets: the NCBI-disease corpus [92] (793 PubMed abstracts annotated with disease mentions) and surgical pathology records (40,443 annotated lung cancer diagnostic records from the Asan Medical Center). Synthetic errors were introduced to mimic real-world typographical patterns. The MLM-based spelling correction achieved F1-scores of 0.72 (NCBI-disease) and 0.73 (surgical pathology records). For NER tasks, spelling correction significantly boosted F1-scores in surgical pathology records from 0.60 to 0.85. This study highlighted the potential of spelling correction models to mitigate data quality issues and improve the accuracy of downstream NLP tasks in clinical settings.
@@
Contextual error correction has seen advancements through the application of LLMs for addressing complex, context-dependent inaccuracies. A highlight in this domain is the MEDIC system introduced by Pais et al. [80], which focuses on preventing medication direction errors in online pharmacies by improving accuracy and standardization during the data entry phase. The study utilized 1.6 million single-line medication directions from Amazon Pharmacy, including raw prescriber directions and pharmacist-verified equivalents. MEDIC employs a three-stage process: pharmalexical normalization, which standardizes and corrects variations in medication terminology and formatting; AI-powered extraction using a fine-tuned DistilBERT model; and semantic assembly with safety guardrails informed by a medication catalog derived from RxNorm, OpenFDA, and Amazon Pharmacy data. Compared to T5-FineTuned and Claude, MEDIC reduced near-miss events by 33 during deployment, with a notable improvement in suggestion adoption rates and a reduction in post-adoption edits.
@@ 4.3 Others
Tavabi et al. [81] highlighted a critical issue in codification errors related to gold standard labels while evaluating TF-IDF, Doc2Vec, and BERT models for assigning current procedural terminology (CPT) codes from operative notes. The study, focused on musculoskeletal procedures, used a dataset of 44,002 operative notes annotated with the 100 most common CPT codes. TF-IDF demonstrated superior performance with an AUROC of 0.96 and accuracy of 0.97, outperforming both Doc2Vec and BERT, likely due to its robustness to data sparsity and noise in clinical notes. Importantly, the study revealed discrepancies in gold standard CPT assignments during experiments, with NLP models flagging potentially mislabeled instances and correcting errors in the provided ground truth in some cases. This underscores the need for refining data labeling processes to enhance the reliability of automated codification systems.
@@ 5 Misinformation
Misinformation correction is less explored compared to detection, with only two studies in this review addressing it [61, 110]. Both studies focused on generating polite, evidence-backed responses using social media posts and academic articles as sources of misinformation and supporting evidence. The two papers either released datasets or built on publicly available datasets, but further datasets are still needed to support advancements in this area.
@@ 5.1.2 Fact-checking and claim verification
Validation and interpretation is the final step in fact-checking and claim verification, focusing on model explainability. This stage aims to clarify the reasoning behind predictions and increase trust and transparency in the model’s conclusions. However, this step is often overlooked, with only a few studies incorporating explanation generation and human evaluation. Kotonya et al. [66] proposed a hybrid explanation generation method using extractive-abstractive summarization, where evidence retrieved by their fact-checking pipeline was summarized into natural language justifications via a BERT-based model. Human annotators then assessed the coherence and relevance of these explanations to ensure they were logically consistent with the claim and its veracity label. In Tan et al. [64], human annotators evaluated claims generated from multiple-choice questions for qualities like fluency, contextuality, and faithfulness. While some studies included expert evaluations for claim quality or evidence matching, few employed specific metrics to measure interpretability. In future studies, effective validation and interpretation methods will be essential for creating more transparent fact-checking systems that can explain decisions comprehensively and reliably.
@@ 6 Hallucination
| Van Veen et al. [132] | general medical | 1 | LLM-generated summaries | Open-i, MIMIC-CXR, MIMIC-III, MeQSum, ProbSum | GPT-4, GPT-3.5, FLAN-T5, FLAN-UL2, Vicuna, Llama-2 | BLEU, ROUGE-L, BERTScore, MEDCON, correctness, completeness, conciseness | automatic, expert |